OpenStudy (zuha):

how to solve: 1/2+1/4*1/4-1/16 A. 11/16 B.9/16 C. 4/16 D.8/16

OpenStudy (uri):

I'll let @GoldPhenoix do it :D

OpenStudy (goldphenoix):

\[\large \large \frac{ 1 }{ 2 }+ \frac{ 1 }{ 4 } * \frac{ 1 }{ 4 } - \frac{ 1 }{ 16 }\] You would want to do Order Of Operation. So what do you do first?

OpenStudy (zuha):

add

OpenStudy (zuha):

and subtract then multiply

OpenStudy (goldphenoix):

Hmm. That's the last thing we do. Do you remember PEMDAS?

OpenStudy (zuha):

no

OpenStudy (zuha):

what is that

OpenStudy (goldphenoix):

PEMDAS is very important! P= Parenthesis E= Exponent M = Multiplication/Division D = Division/Multiplication A = Addition/Subtraction S = Subtraction/Addition. Do we have parenthesis?

OpenStudy (goldphenoix):

We do multiplication first, why? Well because we don't have parenthesis or exponent. So what does (1/4) * (1/4) give us?

OpenStudy (zuha):

1/16

OpenStudy (goldphenoix):

Good job. So now our equation looks like: \[\large \large \frac{ 1 }{ 2 } + \frac{ 1 }{ 16 } - \frac{ 1 }{ 16 }\] Now what do we do next?

OpenStudy (zuha):

add

OpenStudy (zuha):

?

OpenStudy (goldphenoix):

What does (1/2) + (1/16) give us?

OpenStudy (zuha):

is it 1/18

OpenStudy (goldphenoix):

Nope. You want to find a common demoniator. What is the least common multiple of 2 and 16?

OpenStudy (zuha):

4

OpenStudy (goldphenoix):

Hmm. Let me help you out: 2: 2, 4, 6, 8, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 32 16: 16, 32, 48 What number does 2 and 16 share in common?

OpenStudy (zuha):

2

OpenStudy (zuha):

oh

OpenStudy (zuha):

16

OpenStudy (goldphenoix):

Good job.

OpenStudy (goldphenoix):

 What would the numerator be?

OpenStudy (zuha):

dont get it !

OpenStudy (goldphenoix):

We want to add 1/2 and 1/16. Correct? To add them, they both must have the same demoniator. You said that the least common multiple of 2 and 16 is 16. Therefore, 1/16 stay the same because 1/16 already has a demoniator of 16. So if 1/2 has a demoniator of 16, then what is the numerator?

OpenStudy (zuha):

1

OpenStudy (goldphenoix):

Hmm. Well. To get 2 to 16. What do we do?

OpenStudy (zuha):

multiply by 8

OpenStudy (goldphenoix):

So we must do the same thing to the numerator. So what is 1 * 8?

OpenStudy (zuha):

8

OpenStudy (goldphenoix):